OpenStudy (anonymous):

@misty1212 A recipe that makes 6 servings calls for cups of juice. How many cups of juice are there per serving?

OpenStudy (misty1212):

lol

OpenStudy (misty1212):

copy and paste fail on this one

OpenStudy (anonymous):

yep :(

OpenStudy (misty1212):

in any case cups per serving divide the number of cups by the number of servings (6)

OpenStudy (anonymous):

yeah but its trying to use fracions

OpenStudy (misty1212):

what is the number of cups?

OpenStudy (anonymous):

1 1/5

OpenStudy (misty1212):

ok so \[1\tfrac{1}{5}=\frac{6}{5}\] and \[\frac{6}{5}\div6=\frac{1}{5}\]